Warren Company plans to open a new repair service center for one of its electronic products. The center requires an investment in depreciable assets costing $420,000. The assets will be depreciated on a straight-line basis, over four years, and have no expected salvage value. The annual income statement for the center is given below.

Revenues $350,000
Less: Cash operating expenses (140,000)
Depreciation (105,000)
  Income before income taxes $105,000
Less: Income taxes (@40%) 42,000
  Net income $63,000

Required:

1. Using the income approach, calculate the after-tax cash flows.
$

2. Using the decomposition approach, calculate the after-tax cash flows for each item of the income statement and show that the total is the same as the income approach. Enter cash expenses as negative amounts and noncash expenses as positive amounts.

Revenue (after tax) $
Cash expenses (after tax)
Depreciation tax savings
Operating cash flow $

3. What if it is desirable to express the decomposition approach in a spreadsheet format for the four years to facilitate the use of spreadsheet software packages? Express the decomposition approach in a spreadsheet format, with a column for each income item and a total column. Enter the after-tax cash expenses as negative amounts. Enter the after-tax revenues and the tax savings from noncash expenses as positive amounts.

Year Revenue Cash operating
expenses
Noncash operating
expenses
Cash flow
1 $ $ $ $
2
3
4
Answer #1

a)

Using the income approach, the after-tax cash flows =Net Income +Depreciation

                                          = 63000 + 105000

                                            = $ 168000

b)

Revenue (after tax) [350000(1-.40)] 210000
Cash expenses (after tax) [-140000 (1-.40)] - 84000
Depreciation tax savings [105000 * .40] 42000
Operating cash flow 168000

c)

Year Revenue Cash operating
expenses
Non cash operating
expenses
Cash flow
1 350000(1-.40) = 210000 -140000(1-.40)= -84000 105000*.40=42000 168000
2 210000 -84000 42000 168000
3 210000 -84000 42000 168000
4 210000 -84000 42000 168000